OpenStudy (anonymous):

how to integrate when the power is -1 therefore the whole power comes as 0

OpenStudy (anonymous):

You mean\[\int x^{-1}\]

OpenStudy (anonymous):

yeah..:)

OpenStudy (anonymous):

Remember that if \(\large f(x) = \ln(x) \) then \(\large f^\prime(x) = \frac{1}{x}\)

OpenStudy (anonymous):

Yeah, what @wio said.

OpenStudy (anonymous):

Integration is just finding the anti-derivative!

OpenStudy (anonymous):

so if i want to integrate -3x ^ (-1) then what is the answeR?

OpenStudy (anonymous):

\[\int(-3)x^{-1}\,dx=-3\int\frac1x\,dx=-3\ln|x|+C\]

OpenStudy (anonymous):

ok thanks! :)

hartnn (hartnn):

@abs4 \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

Aw dang @hartnn you're like late.

hartnn (hartnn):

lol! late for what?

OpenStudy (anonymous):

The welcome comes first, so that the person who answers the question looks more awesome.

OpenStudy (anonymous):

Like, "Welcome, but we already know the answer you plebe."

hartnn (hartnn):

oh, i didn't want to interrupt the explanation.....